OpenStudy (anonymous):

What is a quick and easy way to work out 26 cubed minus 25 cubed? Thx

OpenStudy (anonymous):

\[26^3-25^3=(26-25)(26^2+25^2+26*25)\]

OpenStudy (whpalmer4):

well, one way would be to use the formula for the difference of cubes: \[a^3 – b^3 = (a – b)(a^2 + ab + b^2)\] In this case, it would be \[(26-25)(26^2+26*25+25^2)=(1)(26^2+26*25+25^2)\]\[ = 676 + 650 + 625 = 1951 \]

OpenStudy (whpalmer4):

is it really any easier? well...

OpenStudy (anonymous):

Thx